# CompilerDream: Learning a Compiler World Model for General Code Optimization

### v2

@@ 3.2 Learning a Compiler World Model
We develop our CompilerDream approach by adapting an advanced model-based RL method, Dreamer (Hafner et al. 2023), as depicted in Figure 2. This approach involves learning a predictive world model $(\hat{p}_{\theta},\hat{r}_{\theta})$ of the compiler environment parameterized by $\theta$ that approximates the underlying transition dynamics $p(o_{t+1}|o_{\leq t},a_{\leq t})$ and reward function $r(o_{\leq t},a_{\leq t})$.
@@
Representation model: $z_{t}\sim q_{\theta}(z_{t}\mid z_{t-1},a_{t-1},o_{t})$ (1)
Transition model: $\hat{z}_{t}\sim p_{\theta}(\hat{z}_{t}\mid z_{t-1},a_{t-1})$
Observation decoder: $\hat{o}_{t}\sim p_{\theta}(\hat{o}_{t}\mid z_{t})$
@@
$\mathcal{L}_{\mathrm{repr}}(\theta)\doteq-\ln p_{\theta}(o_{t}|z_{t})-\ln p_{\theta}(r_{t}|z_{t}).$ (2)
The transition model captures the dynamics of the compiler world model, predicting future neural state ${\hat{z}}_{t}$ directly from $z_{t-1}$ and $a_{t-1}$. A prediction loss minimizes the difference between the estimated state $z_{t}$ and the predicted state ${\hat{z}}_{t}$, simultaneously enhancing the transition model’s accuracy in predicting future states and making the state produced by the representation model easier to predict:
$\mathcal{L}_{\mathrm{pred}}(\theta)\doteq\mathrm{KL}\left[q_{\theta}(z_{t}|z_{t-1},a_{t-1},o_{t})\,\|\,p_{\theta}(\hat{z}_{t}|z_{t-1},a_{t-1})\right].$ (3)
@@
During an optimization sequence, not all passes can effectively transform the code, and the improvements gradually saturate. We employ a reward smoothing technique (Lee, Abbeel, and Lee 2024) to mitigate the sparsity and the long tail distribution of rewards during an episode. This is achieved through the following:
$r_{t}^{\prime}\leftarrow\alpha r_{t-1}^{\prime}+(1-\alpha)r_{t},\;\;t=1,2,\dots$ (4)
with $\alpha\in[0,1)$. Consequently, we train a reward decoder $p_{\theta}(\hat{r}_{t}^{\prime}\mid z_{t})$ to predict the smoothed rewards.
3.3 Learning a Code Optimization Agent
We then train a code optimization agent on top of the world model, which consists of an actor and a critic neural networks parameterized on top of the neural compiler state:
@@ 3.3 Learning a Code Optimization Agent
The critic evaluates the cumulative optimization return that the actor can achieve starting from a given neural state, also known as value or return, denoted as $v_{\xi}(\hat{z}_{t})$. This is trained by minimizing the difference between the predicted value $v_{\xi}(\hat{z}_{t})$ and the Monte-Carlo return or more advanced bootstrapped $\lambda$-return (Sutton and Barto 2018):
@@
$\mathcal{L}_{\text{critic}}(\xi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\left[\sum^{t+H}_{\tau=t}-\log v_{\xi}(V_{\tau}^{\lambda}\mid\hat{z}_{\tau})\right],$ (6)
$V_{\tau}^{\lambda}\doteq\hat{r}_{\tau}+\gamma\begin{cases}(1-\lambda)v_{\xi}(\hat{z}_{\tau+1})+\lambda V_{\tau+1}^{\lambda}&\text{if}\ \tau<t+H\\ v_{\xi}(\hat{z}_{\tau+1})&\text{if}\ \tau=t+H.\end{cases}$
@@
The actor produces a policy $\pi_{\psi}$ that predicts the best optimization pass to choose, represented as an action distribution, which is trained to maximize the simulated return through the REINFORCE policy gradient (Williams 1992):
@@
$(\displaystyle\mathcal{L}_{\text{actor}}(\psi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\Big{[}\sum^{t+H}_{\tau=t}\Big{(}$$-\left(V_{\tau}^{\lambda}-v_{\xi}(\hat{z}_{\tau})\right)\log\pi_{\psi}(\hat{a}_{\tau}|\hat{z}_{\tau})$
$-\eta\,\mathbb{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]\Big{)}\Big{]},$ (7)
where $\mathbb{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]$ is an entropy regularization which encourages exploration, and $\eta$ is a hyperparameter that adjusts the regularization strength.
Both the actor and critic are completely trained on world model-simulated compiler optimization trajectories.
3.4 Data Curation
@@
To facilitate that our CompilerDream method can effectively generalize to unseen situations, a concept known as zero-shot generalization, we have identified three critical factors in preparing our training dataset. First, the dataset should reflect naturalness. During our preliminary experiments, we found that those generated code datasets like Csmith (Yang et al. 2011) and llvm-stress (Lattner and Adve 2004) provide no benefits or even hurt the generalization to real-world scenarios. Second, the dataset should be large to prevent the agent from overfitting to a small number of codes and failing to generalize. Last, the code must exhibit high quality from an optimization perspective. We need data with complex algorithmic logic and potential for optimizations to help the world model and the agent better understand the problem’s intricacies. Datasets like AnghaBench (Da Silva et al. 2021) which consist of millions of human-written codes collected from GitHub, are often too simple to allow for significant optimization improvements.
@@
Therefore, we choose to construct our training datasets on top of the CodeContests dataset released with AlphaCode (Li et al. 2022), which consists of over 13,000 problems of coding competition, and each problem, on average, has hundreds of solutions in multiple languages. We subsample up to ten C++ solutions for each training problem, resulting in 110,240 programs, as our training data, and sample one solution for each of 100 test problems as our validation data.
4 Experiments
We evaluate CompilerDream in three different settings and compare its performance with prior state-of-the-art methods. We first demonstrate that, without considering generalization, CompilerDream excels as a powerful autotuning method, leading the CompilerGym leaderboard (Section 4.2). Then we show that CompilerDream can learn a general compiler world model capable of accurately estimating and selecting pass sequences (Section 4.3) and finally learn a general optimization agent that directly generates a pass sequence tailored to any specific input program. (Section 4.4)
4.1 Evaluation
@@
Our experiments focus on code size reduction, which benefits applications targeting low-resource hardware such as embedded systems. This focus is driven by the practical advantages of code size as a metric: it is cost-effective to construct extensive compilable training and test datasets and to evaluate the optimization performance for code size.
@@
$R_{\mathcal{D}}(\pi)=\left(\prod_{s\in\mathcal{D}}\frac{C(s_{Oz})}{C(s_{t}^{\pi})}\right)^{\frac{1}{|\mathcal{D}|}}$ (8)
where $C$ counts the number of IR-instructions, $s_{Oz}$ is the IR optimized by LLVM -Oz flag, and $s_{t}^{\pi}$ is the final IR produced by the algorithm $\pi$. A value above $1$ indicates that the algorithm performs better than -Oz option on average.
@@
We evaluate our method mainly on benchmarks from the CompilerGym platform (Cummins et al. 2022): benchmark suites including cBench (Fursin et al. 2007), CHStone (Hara et al. 2008), MiBench (Guthaus et al. 2001), and NASA Parallel Benchmarks (NPB) (Bailey et al. 1995), in addition to kernels from open source projects including BLAS (Lawson et al. 1979), Linux, OpenCV (Culjak et al. 2012), and TensorFlow (Abadi et al. 2016). We disregard benchmarks generated by synthetic program generators (Lattner and Adve 2004; Yang et al. 2011) since they do not align with real-world scenarios. We follow the standard data splits of CompilerGym.
Additionally, we evaluate CompilerDream on a large-scale dataset named FormAI (Tihanyi et al. 2023). FormAI comprises a vast collection of AI-generated C programs with diverse functionalities and coding styles. We filtered out codes that failed to compile into CompilerGym benchmarks, resulting in a test set of 109,016 programs.
Inspired by FormAI, we also constructed a dataset of 50 Objective-C programs generated by a Large Language Model to further evaluate CompilerDream’s ability to generalize to different programming languages. All details on benchmarks can be found in Appendix C.
Figure 3: Generel code optimization results: Code size reduction in terms of IR instruction count over LLVM -Oz under different methods. Bars indicate the geometric mean and min-max range across test programs in each benchmark dataset.
4.2 Autotuning: CompilerGym Leaderboard
@@
Table 1: Autotuning results: CompilerDream’s code size reduction and time consumption on cBench dataset compared with other methods from CompilerGym leaderboard (Facebook 2022). Results for methods other than ours are taken directly from the leaderboard, where learning-based methods report only inference time as wall time.
We first use CompilerDream as an autotuning method to verify that its training process has sufficient capability to discover high-quality optimization sequences, supporting our goal of training an effective general code optimization agent.
@@
We target the task of CompilerGym leaderboard (Facebook 2022), optimizing pass sequences across 23 cBench programs. CompilerDream is trained for 25 hours, averaging about one hour per program—a time budget comparable to search methods. After training, besides evaluating single-trial optimization performance (more in Section 4.4), we test CompilerDream+Guided Search, following the leaderboard’s top method. This approach uses the agent’s policy $\pi_{\psi}(a_{t}|s_{t})$ to guide random search and limits search time to 1 minute per program (details in Appendix B.7).

### v3

@@ 3.2. Learning a Compiler World Model
Following the advanced Dreamer approach (Hafner et al., 2023), we build a world model to learn the formulated POMDP process of compiler optimization, as depicted in Figure 2. Concretely, we train a model $(\hat{p}_{\theta},\hat{r}_{\theta})$ of the compiler environment parameterized by $\theta$, which approximates the underlying transition dynamics of optimization process $p(o_{t+1}|o_{\leq t},a_{\leq t})$ and the reward function $r(o_{\leq t},a_{\leq t})$.
@@
(2) Representation model: $z_{t}\sim q_{\theta}(z_{t}\mid z_{t-1},a_{t-1},o_{t}),$
Transition model: $\hat{z}_{t}\sim p_{\theta}(\hat{z}_{t}\mid z_{t-1},a_{t-1}),$
Observation decoder: $\hat{o}_{t}\sim p_{\theta}(\hat{o}_{t}\mid z_{t}),$
@@
(3) $\mathcal{L}_{\mathrm{repr}}(\theta)\doteq-\ln p_{\theta}(o_{t}|z_{t})-\ln p_{\theta}(r_{t}|z_{t}).$
The transition model captures the dynamics of the compiler world model, predicting future neural compiler state ${\hat{z}}_{t}$ directly from $z_{t-1}$ and $a_{t-1}$. A prediction loss minimizes the difference between the estimated neural compiler state $z_{t}$ and the predicted compiler state ${\hat{z}}_{t}$, simultaneously enhancing the transition model’s accuracy in predicting future states and making the neural compiler state produced by the representation model easier to predict:
(4) $\mathcal{L}_{\mathrm{pred}}(\theta)\doteq\mathrm{KL}\left[q_{\theta}(z_{t}|z_{t-1},a_{t-1},o_{t})\,\|\,p_{\theta}(\hat{z}_{t}|z_{t-1},a_{t-1})\right].$
@@
Figure 3. Comparison of the original and smoothed rewards during an episode on a program from the BLAS dataset (Lawson et al., 1979).
[…]
Taking a typical episode as an example: out of 45 actions, only 8 receive non-zero rewards before smoothing, and they are scattered throughout the optimization process. After smoothing, most rewards become non-zero, and abrupt changes between adjacent rewards are eliminated. Moreover, the overall trend remains consistent with the original reward distribution.
@@
We discover that most optimization passes in an optimization sequence do not change the program IR instruction count, leading to sparse rewards. Despite this, these passes are essential as they may modify critical properties, such as instruction order or replacements, that affect the effectiveness of subsequent passes. Furthermore, improvements tend to saturate over time, as large non-zero rewards typically appear early in the sequence and diminish as optimization progresses.
In Figure 3, we present a typical episode of the optimization process: out of 45 actions, only 8 yielded non-zero rewards, whose values decrease as optimization progresses. These properties pose challenges for training the agent effectively, as it receives little to no guidance from most actions, even when some are vital to the final result. Moreover, the world model may struggle to predict reward values accurately, often defaulting to zero. This challenge arises from the significant class imbalance between zero and non-zero rewards, making it difficult to determine both the timing and magnitude of non-zero rewards.
Therefore, we applied reward smoothing to mitigate their sparsity and long-tailed distribution within an episode. This is achieved by adding an exponential decay to rewards:
(5) $r_{t}^{\prime}\leftarrow\alpha r_{t-1}^{\prime}+(1-\alpha)r_{t},\;\;t=1,2,\dots$
with $\alpha\in[0,1)$. Consequently, we train a reward decoder $p_{\theta}(\hat{r}_{t}^{\prime}\mid z_{t})$ to predict the smoothed rewards. As shown in Figure 3, the smoothed rewards provide non-zero feedback for most steps, making them easier for CompilerDream to learn. It is worth noting that the total reward of an episode remains approximately unchanged after smoothing. For all $t\geq 1$, its contribution to the total reward after smoothing approximates the original reward value if the horizon is infinite: $\lim_{n\to\infty}(1-\alpha)\sum_{i=1}^{n}\alpha^{i}r_{t}=r_{t}$. Since most episodes exhibit a long-tailed reward distribution and we use a relatively small smoothing factor ($\alpha=0.6$), the smoothed total reward is expected to closely match the original total reward.
3.3. Learning a Code Optimization Agent
Building on the world model, we implement two agent designs, unifying supervised and reinforcement learning methods for compiler optimization into our world-model-based approaches.
Value prediction agent
This kind of agent adopts a classic approach in compiler optimization, employing a heuristic to guide search methods (Liang et al., 2023; Luk et al., 2009; Stephenson et al., 2003). It comprises a value prediction model $v_{\theta}$ to estimate the effect of an optimization sequence and a search strategy that leverages this model to identify the best sequence.
The value of a program state $s$ is defined as the expected cumulative reward of applying an action sequence to the program. This aligns with the ratio of the reduced IR instruction count achieved between applying action sequence $\{a_{\tau}\}_{\tau=1}^{m}$ and standard $-Oz$ flag, according to the reward function $r$ defined in Eq. (1):
(6) $v(s_{0},\{a_{\tau}\})=\sum_{\tau=1}^{m}r_{\tau}=\sum_{\tau=1}^{m}\frac{C(s_{\tau-1})-C(s_{\tau})}{C(s_{0})-C(s_{b})}=\frac{C(s_{0})-C(s_{m})}{C(s_{0})-C(s_{b})}.$
We leverage our world model to build a value prediction model. As shown Section 3.2, from an initial observation $o_{0}$, and an action sequence $\{a_{\tau}\}$, our world model can simulate an optimization trajectory $\{\hat{z}_{\tau},\hat{a}_{\tau},\hat{r}_{\tau},\hat{o}_{\tau}\}_{\tau=1}^{m}$. Summing up all simulated rewards $\hat{r}_{\tau}$ yields the predicted value:
(7) $v_{\theta}(s_{0},\{a_{\tau}\})=\sum_{\tau=1}^{m}\hat{r}_{\tau}\approx\sum_{\tau=1}^{m}r_{\tau}$
The search strategy then seeks the optimal optimization sequence $\{a_{\tau}\}^{*}$ by evaluating candidates using the value prediction model $v_{\theta}$. Although various search algorithms are applicable, we employ a simple strategy that enumerates action sequences within a fixed search space (detailed in Section 4.3).
Reinforcement learning agent
RL is another popular approach to compiler optimization, learning a policy to select passes sequentially. Unlike prior methods that rely on costly compiler interactions, CompilerDream trains RL agents on world model-simulated compiler optimization trajectories, significantly improving efficiency.
Our RL agent comprises an actor and a critic neural networks, both parameterized based on the neural compiler state:
@@ 3.3. Learning a Code Optimization Agent
Figure 4. Results of general code optimization with RL: Code size reduction in terms of IR instruction count over LLVM -Oz under different methods. Bars indicate the geometric mean and min-max range across test programs in each benchmark dataset.
[…]
The performance of several methods (-O0, Random Policy, -Oz, Random Search, PPO, In-domain CompilerDream, and Zero-shot CompilerDream) on different benchmarks (BLAS, cBench, CHStone, Linux, MiBench, NPB, OpenCV and TensorFlow).
@@
The critic evaluates the $\gamma$-discounted value $v_{\xi}(\hat{z}_{t})$ of simulated neural state $\hat{z}_{t}$ under policy $\pi_{\psi}$. It is trained by minimizing the difference between the predicted value $v_{\xi}(\hat{z}_{t})$ and the Monte-Carlo or more advanced bootstrapped return (Sutton and Barto, 2018), denoted as $V_{t}$:
@@
(9) $\mathcal{L}_{\text{critic}}(\xi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\left[\sum^{t+H}_{\tau=t}-\log v_{\xi}(V_{\tau}\mid\hat{z}_{\tau})\right],$
@@
The actor produces an optimization policy $\pi_{\psi}$ that predicts an action distribution of the best pass to choose, which is trained to maximize the simulated return through the REINFORCE policy gradient (Williams, 1992) with an entropy regularization (Haarnoja et al., 2019):
@@
$(\displaystyle\mathcal{L}_{\text{actor}}(\psi)\doteq\mathbb{E}_{p_{\theta},\pi_{\psi}}\Big{[}\sum^{t+H}_{\tau=t}\Big{(}$$-\left(V_{\tau}-v_{\xi}(\hat{z}_{\tau})\right)\log\pi_{\psi}(\hat{a}_{\tau}\mid\hat{z}_{\tau})$
(10) $-\eta\,\mathbb{H}\left[\pi_{\psi}(\hat{z}_{\tau})\right]\Big{)}\Big{]},$
3.4. Data Curation
@@
To facilitate that our CompilerDream method can effectively generalize to unseen situations, a concept known as zero-shot generalization, we have identified three critical factors in preparing our training dataset. First, the dataset should reflect naturalness. During our preliminary experiments, we found that those generated code datasets like Csmith (Yang et al., 2011) and llvm-stress (Lattner and Adve, 2004) provide no benefits or even hurt the generalization to real-world scenarios. Second, the dataset should be large to prevent the agent from overfitting to a small number of codes and failing to generalize. Last, the code must exhibit high quality from an optimization perspective. We need data with complex algorithmic logic and potential for optimizations to help the world model and the agent better understand the problem’s intricacies. Datasets like AnghaBench (da Silva et al., 2021) which consist of millions of human-written codes collected from GitHub, are often too simple to allow for significant optimization improvements.
Table 2. Comparison of different training datasets for compiler optimization.
Dataset
Name
Number of
IR files
Large-
scale
Human-
written
Code
quality
cBench (Fursin et al., 2007) 23 No Yes High
Mibench (Guthaus et al., 2001) 40 No Yes Low
Csmith (Yang et al., 2011) $\infty$ Yes No High
llvm-stress (Lattner and Adve, 2004) $\infty$ Yes No Low
AnghaBench (da Silva et al., 2021) $1,041,333$ Yes Yes Low
CodeContests (Li et al., 2022) $110,240$ Yes Yes High
@@
Therefore, we choose to construct our training datasets on top of the CodeContests dataset released with AlphaCode (Li et al., 2022), which consists of over 13,000 problems of coding competition, and each problem, on average, has hundreds of solutions in multiple languages. We subsample up to ten C++ solutions for each training problem, resulting in 110,240 programs, as our training data, and sample one solution for each of 100 test problems as our validation data. As shown in Table 2, CodeContests meets all our criteria, whereas other datasets fall short in one or more aspects.
4. Experiments
We conducted extensive experiments across various settings, comparing CompilerDream against state-of-the-art methods. Our study aims to answer the following key research questions (RQs):
  1. RQ 1
    Optimality: Can CompilerDream’s joint training process of the compiler world model and optimization agent discover superior optimization sequences?
  2. RQ 2
    Accuracy: Does CompilerDream’s world model produce accurate simulations of the real optimization process?
  3. RQ 3
    Generalization: Is the policy learned with the world model still effective on various unseen programs?
  4. RQ 4
    Effectiveness: Are all the techniques employed in CompilerDream effective in enhancing optimization results?
In the following sections, we first show that CompilerDream excels as a powerful autotuning method, leading the CompilerGym leaderboard by discovering superior optimization sequences and leveraging accurate world model (Section 4.2), addressing RQ 1 and partially RQ 2. We then demonstrate that its value prediction agent (Section 4.3) and RL agent (Section 4.4) both outperform state-of-the-art methods in various unseen test datasets, addressing RQ 2 and RQ 3. Finally, Section 4.5 presents ablation studies for RQ 4.
4.1. Evaluation
@@
Our experiments focus on code size reduction, benefiting applications on low-resource hardware like embedded systems. This focus stems from the practical advantages of code size as a metric: It is cost-effective to construct compilable training and test datasets and to evaluate the optimization performance for code size.
@@
(11) $R_{\mathcal{D}}(\text{agent})=\left(\prod_{s\in\mathcal{D}}\frac{C(s_{b})}{C(s_{\text{agent}})}\right)^{\frac{1}{|\mathcal{D}|}}$
where $C$ denotes the IR instruction count, $s_{b}$ is the IR state optimized by LLVM’s -Oz flag serving as a baseline, and $s_{\text{agent}}$ is the final IR state produced by agent. A value of $R$ above $1$ indicates superior performance compared to LLVM’s -Oz option.
@@
We evaluate our method mainly on benchmarks from the CompilerGym platform (Cummins et al., 2022): benchmark suites including cBench (Fursin et al., 2007), CHStone (Hara et al., 2008), MiBench (Guthaus et al., 2001), and NASA Parallel Benchmarks (NPB) (Bailey et al., 1995), as well as kernels from open source projects such as BLAS (Lawson et al., 1979), Linux, OpenCV (Culjak et al., 2012), and TensorFlow (Abadi et al., 2016). Synthetic benchmarks from program generators (Lattner and Adve, 2004; Yang et al., 2011) are excluded as they lack real-world relevance. We adhere to the standard data splits of CompilerGym. For benchmarks with a total number of programs more than 100, we use the first 50 programs as the test set, the following 50 programs as the validation set, and all of the rest as the training set. These training and validation sets are only used for in-domain training. The datasets comprising fewer than 100 programs are not applicable for in-domain training; instead, all their programs are allocated to the test set. The number of programs in each dataset after division is detailed in Table 3.
Table 3. Dataset division of 8 CompilerGym benchmarks.
Benchmark Training Split Validation Split Test Split
BLAS 200 50 50
cBench N/A N/A 23
CHStone N/A N/A 12
Linux 13,794 50 50
MiBench N/A N/A 40
NPB 22 50 50
OpenCV 342 50 50
TensorFlow 1,885 50 50
Additionally, we evaluate CompilerDream on a large-scale dataset named FormAI (Tihanyi et al., 2023). FormAI comprises a vast collection of AI-generated C programs with diverse functionalities and coding styles. We filtered out codes that failed to compile into CompilerGym benchmarks, resulting in a test set of 109,016 programs.
Inspired by FormAI, we also constructed a dataset of 50 Objective-C programs generated by a Large Language Model to further evaluate CompilerDream’s ability to generalize to different programming languages. Details can be found in Appendix B.8.
4.2. Autotuning: CompilerGym Leaderboard
Table 4. Autotuning results on cBench, i.e., the CompilerGym leaderboard (Facebook, 2022), where learning-based methods report only inference time following leaderboard convention.
@@ 4.2. Autotuning: CompilerGym Leaderboard
We first validate CompilerDream as an autotuning method to demonstrate its ability to discover high-quality optimization sequences, supporting our goal of achieving superior code optimization.
@@
We target the CompilerGym leaderboard task (Facebook, 2022), optimizing pass sequences for 23 cBench programs. Following the common setup adopted by other methods on the leaderboard, we train CompilerDream using our RL agent on all cBench programs except ghostscript, which is excluded due to its large size that would significantly slow down training. The action space is set to the full action space of all 124 actions in LLVM. CompilerDream is trained for 25 hours, averaging about one hour per program—comparable to the 3600-second wall time of the random search algorithm. Beyond evaluating single-trial optimization performance, we test CompilerDream+Guided Search, inspired by the leaderboard’s leading approach. This method leverages the RL agent’s policy $\pi_{\psi}(a_{t}|s_{t})$ to guide random search, limiting search time to 1 minute per program (details in Appendix B.6).